OpenStudy (anonymous):

Did I do this Algebra question right? 4= [2x+6] I believe the answer is x=-1 and x=5 is this correct? :) Merci beaucoup

OpenStudy (alexwee123):

are [ ] absolute value?

OpenStudy (anonymous):

Est-ce la valeur absolue ou valeur a part entiere (les crochets)

OpenStudy (anonymous):

Yes, yes it is

OpenStudy (anonymous):

ABSOLUTE VALUE

OpenStudy (alexwee123):

then it's wrong for 2nd value of x

OpenStudy (anonymous):

it's -5?

OpenStudy (alexwee123):

ya :)

OpenStudy (anonymous):

Are you lying to me?

OpenStudy (anonymous):

x=-1 and -5

OpenStudy (accessdenied):

\( \lvert 2x + 6 \rvert = 4 \\ 2x + 6 = 4 \text{ and } 2x + 6 = -4 \\ 2x = -2 \text{ and } 2x = -10 \\ x = -1 \text{ and } x = -5 \)

OpenStudy (anonymous):

4=-2x-6 10=-2x -5=x

OpenStudy (alexwee123):